Find x-intercepts and vertex, then sketch graph of f(x)=x2-8x+12
1
Expert's answer
2011-03-29T05:42:46-0400
x-intercepts:
x2-8x +12 = 0
D = 16 - 12 = 4
x1 = 6, x2 = 2.

Vertex: x = -b/2a = 8/2 = 4, y(4) = -4, so the vertex has coordinates (4,-4).